solve the following problem, be sure to show the equation used for the solution.
In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were cast in all, how many did each candidate receive?

Answer by gsmani_iyer(178)


Let the votes received by the loser = x
So the votes received by the winner = x+220 ...(Given)
The total No. of votes cast = 810 ..... (Given)
So x + (x+220) = 810 ........ This equation is important.
So x+x+220 = 810 .... After removing the brackets.
2x + 220 = 810
2x = 810-220 = 590
So x = 590/2 = 295
The votes received by loser = 295
The votes received by the winner = 295+220 = 515 Answer.

I hope this is clear to you now.
gsm